### v1

# Testing the Equivalence Principle on Cosmological Scales Using Peculiar Acceleration Power Spectrum

Abstract
While the (weak) Equivalence Principle (EP) has been rigorously tested within the solar system, its validity on cosmological scales, particularly in the context of dark matter and dark energy, remains uncertain. In this study, we propose a novel method to test EP on cosmological scales by measuring the peculiar acceleration power spectrum of galaxies using the redshift drift technique. We develop an EP estimator, $E_{\rm ep}$, to evaluate the consistency of the peculiar acceleration power spectrum across different tracers. By calculating the ratio of the peculiar acceleration power spectra of tracers, the ensemble average of $E_{\rm ep}$ is expected to be unity if EP holds on cosmological scales for these tracers. We validate this estimator using N-body simulations, focusing on four redshift bins with $z\leq 1.5$ and scales of $k$ in the range of $0.007$ and $0.32$ $h/\rm Mpc$. By measuring $E_{\rm ep}$ using i) different samples of dark matter particle mock data and ii) low-mass and high-mass halo mock data, we find that the measured $E_{\rm ep}$ values are consistent with unity within the $2\sigma$ level, supporting the validity of $E_{\rm ep}$ on the linear cosmological scales. Taking advantage of advanced observing capabilities, such as next-generation facilities that extend beyond the Square Kilometer Array, the proposed method offers a promising approach for future cosmological tests of EP.
methods: data analysis, numerical — cosmology: large-scale structure of Universe, theory
I Introduction
The celebrated general theory of relativity is undoubtedly founded upon the Equivalence Principle (EP), which represents one of its cornerstones. EP states that all non-gravitational phenomena are locally unaffected by gravity, provided that they are performed in a freely falling frame. A specific consequence of this principle is that all objects (including light) fall in accordance with the same laws. Thus far, EP has been subjected to rigorous testing within the solar system (Will, 2014; Touboul et al., 2017). However, its applicability on cosmological scales when considering dark matter and dark energy remains uncertain. In essence, the large-scale structure of the Universe contains valuable information about EP, which can be used to directly evaluate its validity. To date, only indirect methods have been proposed to test EP (Kehagias et al., 2014; Creminelli et al., 2014; Bonvin and Fleury, 2018; Umeh et al., 2021). In this study, we propose a direct test of (weak) EP via the density-weighted peculiar acceleration power spectrum, which can be quantified through the redshift drift technique.
The concept of redshift drift – the change in the redshift of objects following cosmological expansion over time – has been a subject of study for decades. The theory was initially proposed by Sandage (1962); McVittie (1962) and the concept was reinitiated approximately two decades ago by Loeb (1998). Recently, drift motions in perturbed FRW universes, and a scale-dependent dipolar modulation of the acceleration of the expansion rate inferred from observations of objects within the bulk flow have been investigated by Amendola et al. (2008); Tsagas (2010); Colin et al. (2019).
The measurement of redshift drift plays an essential role in real-time cosmology, as it is equivalent to observing the expansion of the Universe in real time. Moreover, the redshift drift is a model-independent method of probing the expansion of the Universe, thereby eliminating assumptions about geometry, clustering, or gravitational behavior. One of the key challenges in measuring redshift drifts arises from the vast difference between the time scales relevant to cosmology and human observation, requiring extremely precise measurements. The primary sources of uncertainty in redshift drift measurements arise from shot noise and redshift measurement precision. Addressing these challenges requires both a high number density of galaxies and spectroscopy with ultra-high spectral resolution.
Recently, there have been conceptual discussions regarding potential measurement techniques and different cosmological observables of redshift drift (Lange and Page, 2007; Uzan et al., 2008a, b; Stebbins, 2012; Neben and Turner, 2013; Kim et al., 2015; Koksbang and Hannestad, 2016; Martins et al., 2016). Although the sensitivity of current facilities is still several orders of magnitude lower than that of the expected signal (Darling, 2012), it is promising to measure redshift drift as future facilities become available (Alves et al., 2019). For example, the Extremely Large Telescope (ELT), with its unique and highly advanced spectrographic capabilities (Liske et al., 2008; Balbi and Quercellini, 2007; Corasaniti et al., 2007; Lake, 2007; Moraes and Polarski, 2011; Geng et al., 2014), will facilitate a significant advancement in our understanding of cosmology by enabling the breaking of degeneracies between cosmological parameters. As recently observed by Klöckner et al. (2015); Kang et al. (2025), low redshift measurements are, in principle, feasible with the Square Kilometre Array (SKA). Similarly,  Yu et al. (2014) proposed that redshift drift measurements at intermediate redshifts could be conducted using intensity mapping techniques (Peterson et al., 2009; Chen, 2012), such as those employed by the Canadian Hydrogen Intensity Mapping Experiment (CHIME) (Bandura et al., 2014). This analysis (Neben and Turner, 2013) should also be applicable to the Hydrogen Intensity and Real-time Analysis eXperiment (HIRAX).
To rigorously test whether EP holds on cosmological scales, it is essential to examine the statistical internal consistency of the peculiar acceleration power spectrum of galaxies. This is a logical consequence of the EP test. As a first step, we propose an EP estimator to evaluate the principle and perform tests on two different mock data sets: dark matter (DM) particles and DM halos. In this study, we explore the potential of radio observations, focusing on an HI galaxy survey beyond the capabilities of the SKA. This survey is expected to cover $10^{8}$ HI galaxies at $z\leq 1.5$ Klöckner et al. (2015); Kang et al. (2025). Moreover, we make a forward-looking assumption that sufficiently long integration times will enable redshift measurements with a precision of $\sim 1~{}{\rm cm/s}$. This precision is significantly higher than the expected capabilities of the SKA redshift drift experiment (Klöckner et al., 2015).
In the following, we will present a theoretical methodology for statistical measurements of the peculiar acceleration of galaxies, which will then be employed to quantify the EP test using mock data.
II Methodology
The galaxy acceleration along the line of sight (LOS), analogous to galaxy velocity, can be decomposed into two distinct components: the background cosmic acceleration, $a_{\rm cos}$, and a perturbed component, namely the peculiar acceleration, $a_{\rm pec,\parallel}$. This decomposition can be expressed as
$a_{\|}\equiv c\dot{z}=a_{\rm cos}+a_{\rm pec,\parallel}\,,$ (1)
where $c$ is the speed of light. As initially demonstrated by (Sandage, 1962), in a homogeneous and isotropic spacetime, the first term is directly linked to the Hubble function through the relation:
$a_{\rm cos}=c(1+z)H_{0}-cH(z)\,,$ (2)
where $H_{0}$ represents the present-day value of the Hubble parameter, $H(z)$. The second term of Eq. 1 accounts for the effect of local gravitational interactions, which cause the object to accelerate relative to the Hubble flow. This effect is determined by the gradient of the gravitational potential ($\phi$):
$a_{\rm pec,\parallel}(\bm{x},z)\equiv-\frac{\nabla\phi}{a}\cdot\hat{z}\,,$ (3)
where the unit vector $\hat{z}$ represents the LOS direction, and the gradient is computed in comoving coordinates.
Note that when applying the redshift drift technique to detect peculiar accelerations, the full expression for the time derivative of galaxy redshift, up to first order in metric perturbations and in $v/c$. In writing Eq. 3, we have made the following simplifications: i) we neglect all perturbative terms related to the local observer (us), as these terms vanish in the power spectrum estimator proposed in this study; ii) we omit two source-related terms: one analogous to the integrated Sachs-Wolfe (ISW) effect and another proportional to the peculiar velocity $v_{\rm pec}$. These terms are subdominant compared to the peculiar acceleration term, as demonstrated in Uzan et al. (2008b). Furthermore, our analysis is conducted in the real space under the plane-parallel limit approximation, though similar conclusions can be extended to the redshift space or angular statistics at linear scales.
II.1 Estimator
To test EP, we propose measuring the density-weighted peculiar acceleration power spectrum of galaxies using the redshift drift method. The corresponding acceleration power spectrum, $P_{a_{\parallel}a_{\parallel}}$, is expressed as follows,
$(2\pi)^{3}\delta^{D}(\bm{k}-\bm{k}^{\prime})P_{a_{\parallel}a_{\parallel}}(\bm{k})$ (4)
$=$ $(1+\delta_{g})a_{\parallel}(1+\delta_{g}^{\prime})a_{\parallel}^{\prime}\right>$
$=$ $(1+\delta_{g})a_{{\rm pec},\parallel}(1+\delta_{g}^{\prime})a_{{\rm pec},\parallel}^{\prime}\right>$
$\approx$ $a_{{\rm pec},\parallel}a_{{\rm pec},\parallel}^{\prime}\right>\,.$
where, in the second equality, the constant term $a_{\rm cos}$ contributes only to the $k=0$ mode and is therefore ignored, and $^{\prime}$ means that we ignore the Dirac-delta function in the power spectrum definition. This is because the $k=0$ mode represents the spatially uniform component of the field, which does not carry information about the spatial structure or clustering of galaxies. In the last equality, we restrict our analysis to linear scales and neglect the second-order term $\left\langle\delta_{g}a_{\rm pec,\|}\right\rangle\approx b_{g}\left\langle\delta_{m}a_{\rm pec,\|}\right\rangle$, where $\delta_{g}$ and $\delta_{m}$ are the galaxy and matter density contrasts, respectively, and $b_{g}$ is the galaxy bias. The feasibility of this approximation is validated in Appendix A, where we demonstrate that the second-order term is negligible on linear scales and does not significantly impact our results.
Consequently, $P_{a_{\parallel}a_{\parallel}}$ is unaffected by the galaxy bias at linear scales. Based on this, we construct the following estimator, denoted as $E_{\rm ep}$, to test EP:
$E_{\rm ep}(\bm{k})=\frac{P^{i}_{a_{\parallel}a_{\parallel}}(\bm{k})}{P^{j}_{a_{\parallel}a_{\parallel}}(\bm{k})}\,.$ (5)
Here, the indices $i$ and $j$ represent distinct sets of a given acceleration tracer. The primary advantage of this estimator lies in its ability to eliminate cosmic variance through the use of a ratio. By constructing the estimator as a ratio of two correlated quantities, the fluctuations contributed by cosmic variance exactly cancel out. This cancellation enables a more precise and reliable test of EP, significantly enhancing the robustness of the results.
To account for the effects of shot noise, we subtract the shot noise term $\langle a_{{\rm pec},\parallel}^{2}\rangle/n$ from the measured power spectrum, where $n$ represents the number density of a tracer (halos or particles in this study). This subtraction removes discrete sampling effects, yielding the corrected power spectrum:
$(2\pi)^{3}\delta^{D}(\bm{k}-\bm{k}^{\prime})P_{a_{\parallel}a_{\parallel}}(\bm{k})=\left<a_{{\rm pec},\parallel}a_{{\rm pec},\parallel}^{\prime}\right>-\frac{\left<a_{{\rm pec},\parallel}^{2}\right>}{n_{h}}\,,$ (6)
where $a_{{\rm pec},\parallel}^{2}\right>$ is the dispersion of $a_{{\rm pec},\parallel}$ for such tracer.
If EP holds on cosmological scales for such tracers, the ensemble average of $E_{\rm ep}$ is expected to be unity. This expectation arises because the EP implies that all tracers should respond to gravitational forces in the same way, leading to a consistent relationship between their acceleration fields. Deviations from unity in the ensemble average of $E_{\rm ep}$ would indicate a violation of the EP or the presence of systematic effects.
II.2 Simulation
To validate EP using the estimator $E_{\rm ep}$, we conducted two distinct simulation tests: one using DM particles as tracers of acceleration, and the other using DM halos as tracers. In the first test, we introduced random redshift measurement errors to each DM particle. Given the large number of particles in the simulation, the impact of these errors on the estimator can be well quantified and does not significantly bias the results. In the second test, we assumed an ideal scenario with no redshift errors due to the relatively low number of halos in our simulation. Introducing redshift errors to individual halos would significantly increase uncertainties in the measured acceleration power spectrum, degrading the quality of the EP test.
For both simulation tests, we performed an N-body simulation using the Gadget4 (Springel et al., 2021) code, adopting a $\Lambda$CDM cosmology with parameters: $\sigma_{8}=0.8228$, $n_{s}=0.96$, $\Omega_{m}=0.307115$, $\Omega_{b}=0.049$, and $h=0.68$, compatible with the Planck 2018 measurements within the $2\sigma$ level (Planck Collaboration et al., 2020). The simulation employs a box size of $L_{\rm box}=1024~{}{\rm Mpc}/h$ and a total of $N_{\rm p}=1024^{3}$ particles. Moreover, we employ the Cloud-in-Cell (CIC) method to map the particle/halo acceleration field onto a regular $1024^{3}$ grid.
II.2.1 DM mock generation
To simulate measurement errors in redshift, reflecting a more realistic observational scenario, we introduced redshift errors in each particle. Specifically:
1) For each redshift snapshot, randomly select $10^{8}$ particles from the full set of $1024^{3}$ DM particles in the simulation. This subset is labeled as “group1”. Then, exclude the particles in “group1” from the full set. From the remaining particles, randomly select another $10^{8}$ particles to form “group2”.
2) Adding random Gaussian redshift noise independently to each particle in both groups. The noise was generated with a mean of zero and a standard deviation of $\sigma_{z_{\parallel}}=f\sigma_{\rm SKA}$, where $\sigma_{\rm SKA}=10~{}{\rm m/s}$ (Klöckner et al., 2015) represents the typical redshift measurement uncertainty expected from SKA observations (an integration time with $\mathcal{O}(1)-\mathcal{O}(10)$ hours per pointing). Here the parameter $f$ is a scaling factor applied to $\sigma_{\rm SKA}$. In this study, we consider two noise levels for enhancing the SKA observation accuracy: Case A with $f=0.001$ and Case B, with $f=\sqrt{2}\times 0.001$, which corresponds to a higher noise level, doubling the noise in the power spectrum.
3) For a given redshift, the acceleration power spectra of the two groups are calculated using Eq. 6, with corrections applied to account for systematic effects. First, we subtract the power spectrum of Poisson shot noise. Additionally, we subtract the mean power spectrum of the random redshift noise, which arises due to the presence of Gaussian-distributed accelerations. This mean noise power spectrum is estimated by computing the power spectrum of $10^{8}$ randomly distributed particles, each assigned Gaussian-distributed accelerations corresponding to either Case A or Case B. The amplitude of this noise power spectrum is approximately $\sim 10^{-16}({\rm Mpc}/h)^{3}{\rm cm^{2}/s^{4}}$.
4) The covariance matrix for the acceleration power spectrum is accurately estimated using the jackknife method (Norberg et al., 2009) on a single mock dataset. Jackknife realizations are constructed by sequentially deleting one sample at a time and calculating the power spectrum from the remaining data. Specifically, using the delete-one jackknife approach, the simulation box is divided into $n_{s}=64$ equal-volume sub-boxes. Each jackknife box contains $16^{3}$ cells on the CIC mesh. The covariance matrix is estimated as
$C_{ij}=\frac{n_{s}-1}{n_{s}}$ $\sum_{l=1}^{n_{s}}$ $\left[E_{\rm ep}^{[l]}\left(k_{i}\right)-\langle E_{\rm ep}\left(k_{i}\right)\rangle\right]$ (7)
$\times$ $\left[E_{\rm ep}^{[l]}\left(k_{j}\right)-\langle E_{\rm ep}\left(k_{j}\right)\rangle\right]\,,$
where $E^{[l]}_{\rm ep}$ is the estimate of the $l$-th realization, and $\langle E_{\rm ep}\rangle$ is the mean value from $n_{s}$ jackknife realizations. Analysis of the mock data reveals that the resulting covariance matrix is almost diagonal, with negligible correlations between different $k$-bins.
II.2.2 Halo mock generation
$z$ $n_{h}$ (low/high) $M_{h}\times 10^{12}~{}M_{\odot}/h$ (low/high)
1.5 $360821/1082464$ $[2.73,3.41]/[3.41,847.050]$
1.0 $466526/1399581$ $[2.73,3.49]/[3.49,1387.33]$
0.5 $536364/1609094$ $[2.73,3.58]/[3.58,2421.35]$
0.0 $550123/1650372$ $[2.73,3.66]/[3.66,5085.91]$
Table 1: Number of halos in the “low-mass” and “high-mass” groups along with their associated mass ranges at different redshifts.
In addition to testing EP using DM particle simulation data, we also conduct the second test by comparing the acceleration power spectra of DM halos with different masses. In the simulation, DM halos are identified using the group finder ROCKSTAR (Behroozi et al., 2013). We selected all halos containing at least 10 DM particles, corresponding to $M_{h}\geq 2.73\times 10^{12}~{}M_{\odot}/h$. These halos are subsequently divided into two mass bins, referred to as “low-mass” and “high-mass”, with the former containing halos of lower mass and the latter containing halos of higher mass. For each fixed redshift, the ratio of the number of low-mass halos to high-mass halos is maintained at approximately $1/3$. This ratio ensures that the shot noise contributions from the small and large mass bins are nearly equal, as confirmed by our measurements. The detailed halo numbers and masses are summarized in Tab. 1. The covariance matrix for $E_{\rm ep}(k)$ is estimated using the same approach described in the previously detailed jackknife method.
III Results
Figure 1: Measurement of $E_{\rm ep}$ for DM particles for testing the Equivalence Principle. The ratio of acceleration power spectra for “group1” and “group2”, each containing $10^{8}$ particles, and the Gaussian redshift error were additionally added in each particle, with zero mean and standard deviation $\sigma_{z_{\parallel}}$, where $f=0.001$ (left) and $f=0.001\times\sqrt{2}$ (right). Four redshifts are considered, and the $k$ values are in the range of $k\in[0.007,0.32]$, with $10$ bins divided on the logarithmic scale. The black solid line indicates the expectation $E_{\rm ep}=1$ if the Equivalence Principle exactly holds. The error bars correspond to the $2\sigma$ level, estimated from the jackknife scheme applied to the simulation box. As seen, in most cases, the $E_{\rm ep}$ values are consistent with unity at the $2\sigma$ level. Note that the bin centers have been slightly adjusted for clarity in the illustration.
$z$ (particle) $\delta_{\rm ep}$ (“A”/“B”) $\sigma_{\delta}$ (“A”/“B”) $\chi^{2}$ (“A”/“B”)
0.0 $0.0004$/$-0.0087$ 0.0098/0.0162 2.71/6.03
0.5 $-0.0097$/$-0.0069$ 0.0081/0.0127 9.17/8.69
1.0 $-0.0040/0.0008$ 0.0064/0.0096 4.05/9.17
1.5 $-0.0018/0.0170$ 0.0057/0.0083 4.50/13.54
Table 2: Best-fit values of $\delta_{\rm ep}$ and the corresponding $1\sigma$ statistical uncertainties, along with the resulting $\chi^{2}$ values for four $z$-bins, are presented for the DM particle simulation data. The results compare two cases (“A” and “B”) with varying SKA observation accuracies.
Figure 2: Same as in Fig. 1, but for the halo mock data, calculated by taking the ratio of acceleration power spectra for the “high-mass” and “low-mass” halo catalogs. In most cases, the derived $E_{\rm ep}$ is consistent with unity within the $2\sigma$ uncertainty level.
$z$ (halo) $\delta_{\rm ep}$ $\sigma_{\delta}$ $\chi^{2}$
0.0 0.0510.051~{}~{}0.0510.051 0.054 4.31
0.5 $-0.021$ 0.038 4.91
1.0 $-0.042$ 0.040 4.09
1.5 $-0.068$ 0.044 12.09
Table 3: Best-fit values of $\delta_{\rm ep}$ and the corresponding $1\sigma$ statistical uncertainties, along with the resulting $\chi^{2}$ values for four $z$-bins, are presented for the DM halo datasets.
Here, we present the measured $E_{\rm ep}$ values for DM particle and halo mock data across four redshift bins ($z=0.0,0.5,1.0,1.5$), with 10 $k$-bins ranging from $0.007$ to $0.32~{}h/\rm Mpc$. The $k$-bins are evenly spaced on a logarithmic scale. For $k\gtrsim 0.3~{}h/\rm Mpc$, the measured power spectra begin to deviate significantly from the linear prediction due to the non-linearity of structure growth (see Appendix A for details). As a result, we restrict our analysis to scales below this threshold.
To quantitatively characterize the degree of deviation from the expected value, we propose a simple one-parameter model to fit the observed data $E^{\rm obs}_{\rm ep}(k)$ over $k$ bins. The parameterization is expressed as:
$E^{\rm model}_{\rm ep}=1+\delta_{\rm ep}\,,$ (8)
where $\delta_{\rm ep}$ represents the deviation from unity, which corresponds to the expected value if EP is perfectly satisfied. The corresponding $\chi^{2}$ is given by
$\chi^{2}=\sum_{k}\frac{\left[E^{\rm obs}_{\rm ep}(k)-E^{\rm model}_{\rm ep}\right]^{2}}{\sigma^{2}_{k}}\,,$ (9)
where correlations between different bins are neglected, as they are negligibly small. The variance $\sigma^{2}_{k}$ in the $k$-th bin is estimated using the jackknife method based on the simulations. Considering that one parameter $\delta$ and 10 $k$-bins are used in the fitting, the degrees of freedom thus are $N_{\rm dof}=10-1=9$.
Fig. 1 presents the measured $E_{\rm ep}$ for DM particle mock data, which serves as a test of the Equivalence Principle. Specifically, we compute the ratio of acceleration power spectra for two groups (as described in Sect. II.2.1). Two different levels of redshift measurement errors are introduced in the simulation particles: the left and right panels correspond to $f=0.001$ and $f=0.001\times\sqrt{2}$, respectively. As shown, the error bars in the right panel are visibly larger than those in the left panel, which is due to the relatively larger $f$ value.
Based on the measured $E_{\rm ep}$, the model fitting results are summarized in Tab. 2, presenting the best-fit values of $\delta_{\rm ep}$, their associated $1\sigma$ statistical uncertainties, and the resulting $\chi^{2}$ values for the four $z$-bins. As seen, the average error in the larger $f$ case is greater than that in the smaller $f$ case by a factor of $1.56$, which is almost consistent with the expected scaling of $\sqrt{2}$. This result demonstrates that the errors are significantly influenced by the redshift measurement error. More importantly, the $E_{\rm ep}$ values are generally consistent with unity within the $2\sigma$ level across most cases. As $N_{\rm dof}=9$, all $\chi^{2}$ values fall within the range of $\chi^{2}\in[0.51,12.48]$, consistent with the $\chi^{2}$ distribution at the $2\sigma$ confidence level. This result validates the $E_{\rm ep}$ estimator as a reliable tool for testing EP at the linear cosmological scales of $k\in[0.007,0.32]~{}h/\rm Mpc$.
In Fig. 2, we present the measured $E_{\rm ep}$ values for the DM halo mock data across four redshift bins. As shown, the $E_{\rm ep}$ values generally agree with the expected values within the $2\sigma$ level. The error bars increase gradually with larger $k$ values, primarily due to the growing influence of shotnoise at smaller scales. After fitting the model, we find that the best-fit $\delta_{\rm ep}$ values are consistent with zero within the $2\sigma$ level, indicating that deviations from unity are not statistically significant. Furthermore, the $\chi^{2}$ values fall within the expected range, supporting the validity of the proposed $E_{\rm ep}$ estimator for the EP test based on the halo simulation data.
IV Conclusion
The (weak) Equivalence Principle (EP) has been extensively tested within the Solar System; however, its validity on cosmological scales remains an open question. In this study, we propose a novel approach to test EP on cosmological scales by measuring the peculiar acceleration power spectrum of galaxies through the redshift drift technique.
To achieve this, we develop an EP estimator, $E_{\rm ep}$, designed to evaluate the consistency of peculiar acceleration power spectra across different tracers. If EP holds on cosmological scales for these tracers, the ensemble average of $E_{\rm ep}$ is expected to equal one. As a first step toward testing EP on cosmological scales, we validate this estimator using N-body simulations. Applying this approach, we measure $E_{\rm ep}$ at four redshift bins ($z=[0,0.5,1,1.5]$) on the linear scales of $k\in[0.007,0.32]~{}h/\rm Mpc$. This analysis is conducted using two distinct tests: i) DM particle mock data, incorporating redshift measurement errors, and ii) mock data for low-mass and high-mass DM halos.
To quantitatively assess deviations of the measured $E_{\rm ep}$ from the expected value of unity, we introduce a simple one-parameter model to fit the observed $E_{\rm ep}$. The results indicate that the deviations in different cases approach zero within the $2\sigma$ level, confirming the validity of $E_{\rm ep}$ on the linear cosmological scales.
This study serves as a proof of concept for testing EP using the N-body simulation. Since the simulation do not include any sources that could introduce EP violations, our estimates are expected to confirm the validity of $E_{\rm ep}$ on linear scales, which is indeed verified by our results. In future study, we will apply the $E_{\rm ep}$ estimator to real, high-precision observational data. Additionally, we will incorporate potential EP violations into the simulations to evaluate how effectively the estimator detects such deviations and assess its robustness in analyzing such simulated datasets.
Acknowledgement: This work is supported by National SKA Program of China (2020SKA0110401, 2020SKA0110402, 2020SKA0110100), the National Key R&D Program of China (2020YFC2201600, 2018YFA0404504, 2018YFA0404601), the National Science Foundation of China (12203107, 12473097, 12073088, 12373005, 12273121), the China Manned Space Project with No. CMS-CSST-2021 (A02, A03, B01), the Guangdong Basic and Applied Basic Research Foundation (2024A1515012309). We also wish to acknowledge the Beijing Super Cloud Center (BSCC) and Beijing Beilong Super Cloud Computing Co., Ltd (http://www.blsc.cn/) for providing HPC resources that have significantly contributed to the research results presented in this paper.
Appendix A Validation of acceleration power spectrum
Figure 3: Comparison of halo acceleration power spectra from simulations and linear theory predictions for four redshift snapshots: low-mass halos (left) and high-mass halos (right).
The power spectra of the observed density field, $P_{\delta\delta}$, and the peculiar velocity field, $P_{vv}$, are related to the matter power spectrum, $P_{m}$, through the following relations:
$P_{\delta\delta}=b^{2}P_{m}\quad\text{ and }\quad P_{vv}=\left(\frac{aHf}{k}\right)^{2}P_{m}\,,$ (10)
where $a$ is the dimensionless scale factor, $H$ is the Hubble parameter, and $b$ is the linear bias factor. In a standard matter-dominated cosmology, the dimensionless growth rate $f$ is an analytical function of the matter density parameter $\Omega_{m}$, with a weak dependence on the cosmological constant $\Omega_{\Lambda}$. Similarly, within the framework of linear cosmological theory, the acceleration power spectrum is directly linked to the matter power spectrum.
As shown in Eq. 3, the peculiar acceleration represents the integrated effect of excess gravitational forces caused by density fluctuations across the observable universe. Given a field of density perturbations, these accelerations can be directly computed using the Poisson equation. In comoving coordinates, the Poisson equation for the perturbed potential $\phi$ in a matter-dominated universe is:
$\nabla^{2}\phi=\frac{3}{2}\Omega_{m}H^{2}a^{2}\delta_{m}\,,$ (11)
where $\delta_{m}$ is the matter density contrast. Transforming it into Fourier space yields:
$-k^{2}\phi(\bm{k},z)=\frac{3}{2}\Omega_{m}H^{2}a^{2}\delta_{m}\,.$ (12)
From this, the linear relationship between the matter density contrast $\delta_{m}$ and the gravitational acceleration $\bm{a}_{\rm pec}$ can be derived. Therefore, the power spectrum of the peculiar gravitational acceleration is:
$P_{aa}(k,z)\equiv\left\langle\bm{a}_{\rm pec}\cdot\bm{a}_{\rm pec}^{*}\right\rangle=\left(\frac{3aH^{2}\Omega_{m}}{2k}\right)^{2}P_{m}(k,z)\,.$ (13)
The power spectrum of the peculiar acceleration along LOS is then:
$P_{a_{\parallel}a_{\parallel}}(k,z)$ $=$ $\frac{1}{3}P_{aa}(k,z)=\frac{3}{4}\left(\frac{aH^{2}\Omega_{m}}{k}\right)^{2}P_{m}(k,z)$ (14)
$=$ $\frac{3}{4}\left(\frac{H^{2}_{0}\Omega_{m0}}{a^{2}k^{2}}\right)^{2}P_{m}(k,z)\,.$
In Fig. 3, we compare the power spectrum $P_{a_{\|}a_{\|}}$ with the corresponding measurements derived from low- and high-mass halo catalogs at four distinct redshifts. The shot noise contribution has been subtracted from the measured spectra. On linear scales ($k\lesssim 0.06~{}h/\rm Mpc$), the measurements exhibit excellent agreement with the predictions of linear theory for both halo mass bins. This agreement supports the earlier argument in the text that higher-order terms, such as $\delta_{g}a_{\mathrm{pec},\|}$, can be neglected on linear scales. At smaller scales, however, the power spectra deviate from the linear predictions due to the influence of non-linear evolution and halo assembly processes.
We also observe that, on scales $k\lesssim 0.3~{}h/\rm Mpc$, the power spectra for the two halo mass bins at the same redshift snapshot are nearly identical. This indicates that $E_{\rm ep}$ remains robust across this scale range. A more detailed discussion of this result is provided in Sect. III.